# Story-Adapter: A Training-free Iterative Framework for Long Story Visualization

### v1

@@ 4 Experiments
In this section, we first introduce the datasets, the evaluation metrics, and implementation details. Then we compare Story-Adapter with previous AR-based and RI-based methods for visualization of both regular-length and long stories. Finally, we validate the effectiveness of the proposed iterative paradigm and Global Reference Cross-Attention (GRCA) through extensive ablations. Additional experimental results, comparison on subject-consistent generation, and human evaluation can be found in the Appendix.
4.1 Dataset and Evaluation
We use the StorySalon dataset (Liu et al., 2024) to benchmark performance for regular-length story visualization. For long story visualization, we curate multiple long stories using GPT-4o (OpenAI, 2024). To evaluate the efficacy of Story-Adapter, we report CLIP text-image similarity (CLIP-T) (Radford et al., 2021), average Fréchet Inception Distance (aFID) (Cheng et al., 2024), and Character-Character Similarity (aCCS) (Cheng et al., 2024). CLIP-T is to measure image-text alignment, both aFID and aCCS are used to evaluate semantic consistency among generated images.
4.2 Implementation Details
To ensure a fair comparison, we used the weights of IP-Adapter (Ye et al., 2023) and IP-AdapterXL (Ye et al., 2023), respectively, resulting in two models: Story-Adapter and Story-AdapterXL. We utilized DDIM (Song et al., 2020a) for 50-step sampling with an unclassified classifier guidance score set to 7.5. For the hyperparameters in our iterative paradigm, we set the number of story iterations to 10 by default. The weight factor $\lambda$ is set to 0.3 for the initial iteration and 0.5 for the final iteration, with linearly interpolated values for the intermediate iterations by our linear weighting strategy.
4.3 Regular-length Story Visualization
Based on the standard setup on StorySalon dataset (Liu et al., 2024), we compare with existing story visualization methods and Stable Diffusion Model (SDM) baselines, including StoryDiffusion (Zhou et al., 2024), StoryGen (Liu et al., 2024), AR-LDM (Pan et al., 2024), SDM (Rombach et al., 2022), Finetuned-SDM (fine-tuned on StorySalon), and Prompt-SDM. For Prompt-SDM, we use prompts of “cartoon-style images”. To adhere to copyright restrictions and ensure fair comparisons, we exclusively utilize text prompts from the open-source subset of the StorySalon test set for evaluation. This subset comprises 6,026 prompts, with an average of 14 frames per story and the longest story containing up to 44 frames.
Table 1: Quantitative comparison for regular-length story visualization.
@@ 4.3 Regular-length Story Visualization
Story-Adapter (Ours) 0.305 0.760 16.52
@@
Story-AdapterXL (Ours) 0.310 0.818 14.63
@@
Story-Adapter (Ours) 0.307 0.754 98.51
@@
Story-AdapterXL (Ours) 0.318 0.802 94.30
@@
CLIP-T results in Tab. 2 show that Story-Adapter and StoryDiffusion (Zhou et al., 2024) visualize content more aligned to the text prompt than previous story visualization models (AR-LDM and StoryGen). Meanwhile, since neither Story-Adapter nor most baselines are trained on the StorySalon dataset, we introduce aFID and aCCS metrics for a fair evaluation of the character consistency among generated story images. Results of aFID and aCCS in Tab. 2 illustrate that Story-Adapter achieves higher semantic consistency of the generated images compared to StoryDiffusion. Such results validate the effectiveness of our design for coherent image sequence visualization.
@@
In Fig. 4, we provide the qualitative comparison results of the open-ended story visualization. Although AR-LDM and StoryGen generate coherent image sequences based on story prompts, the quality of the generated images degrades when story length increases due to the error accumulation issue of the AR paradigm. Results of StoryDiffusion (Zhou et al., 2024) and Story-Adapter show satisfactory story visualization performance. However, StoryDiffusion cannot maintain consistency between certain subjects due to lacking global story comprehension (e.g., “cat” in Fig. 4). Additionally, since StoryDiffusion requires the first few generated images as references, the visualization results are affected by the reference image flaws (e.g., “closed-eye issue” in Fig. 4). In comparison, Story-Adapter performs better in regular-length story visualization benefited from the global features engaged in GRCA.
Figure 5: Qualitative comparisons for long story visualization. The image sequences in orange and blue boxes are generated by StoryDiffusion and Story-Adapter, respectively. Story-Adapter shows advantages in generating semantic consistency and character interactions. Zoom in for a better view.
Figure 6: Ablation study of iterative paradigm: the effect of the iterative paradigm and the impact of different fixing $\lambda$. Zoom in for a better view. See Appendix for results with more iterations.
@@ 4.4 Long Story Visualization
To better evaluate generative quality for long story visualization (i.e., up to 100 frames), we compare to subject-consistent image generation model IP-Adapter (Ye et al., 2023) in addition to existing story visualization methods. SDM baselines are not included in comparison as they are not suitable to generate long consistent content. We use GPT-4o (OpenAI, 2024) to generate 20 long story cases of ten 50-sentence descriptions and ten 100-sentence descriptions.
@@
The quantitative results in Tab. 2 show that our Story-Adapter significantly improves the semantic consistency and the generative coherence for fine-grained interactions for long story visualization compared to existing models. Notably, IP-Adapter employs the same guidance image that leads to less aFID. In contrast, our method improves visual consistency without the need to fix the same reference image.
@@
Fig. 5 shows the visualization results for long stories, indicating that Story-Adapter can generate high-quality, thematically consistent long image sequences based on the text prompts. In particular, StoryDiffusion cannot convey interactions between multiple characters correctly (e.g., “turtle lifting the fishbone trophy” in the 34-th frame and “rabbit running past the camel” in the 46-th frame), whereas Story-Adapter visualizes the interactions between the characters accurately while maintaining subject consistency.
Computational Cost Comparison.
We evaluate the computational cost of single-image generation using CSA in StoryDiffusion (Zhou et al., 2024) and the proposed GRCA with varying numbers of reference images, under the base attention setting for fair comparison. FLOPs are calculated within the diffusion model UNet. As shown in Fig. 8, as the number of reference images increases, StoryDiffusion experiences a significant rise in computation in terms of FLOPs, while Story-Adapter and Story-AdapterXL are slightly affected. This demonstrates the potential of modeling on global embeddings as in GRCA to efficiently sustain global story semantics for long story visualization.
@@
Table 3: Quantitative ablation studies of the design choices of Story-Adapter.
@@
Iteration Paradigm, $\lambda$ = 0.3 0.320 0.760 101.55
Iteration Paradigm, $\lambda$ = 0.5 0.261 0.753 81.72
GRCA 0.322 0.757 105.17
CSA 0.315 0.768 102.44
@@
Figure 8: Computational cost of single image generation under different number reference images.
4.5 Ablation Study
Global Reference Cross-Attention.
We ablate the effect of global semantics modeling by GRCA for long story visualization. Specifically, for each image visualization in the sequence, we only use the single reference image at the corresponding index during the iteration as guidance. By establishing a global comprehension of the story for the diffusion model, Story-Adapter maintains the semantic consistency in the generated image sequence (Tab. 3 and Fig. 4.4).
@@ 4.5 Ablation Study
We conduct ablation experiments to evaluate the effect of the proposed iterative paradigm for long story visualization and to validate our linear weighting strategy compared to the fixed weight factors. As shown in Tab. 3 and Fig. 6, the iterative paradigm improves generation quality for fine-grained interactions and semantic consistency. This is mainly because the iterative paradigm offers a global view of the entire story, thus reducing error accumulation and alleviating the propagation of the reference image flaws. A fixed weight factor of 0.3 minimally impacts visualization during iteration, while a fixed factor of 0.5 leads to excessive consistency in the image sequence. This enables flexibility within the iterative paradigm.
@@
To ablate the effect of the proposed initialization, we use a sequence of images consisting of the characters as reference images (i.e., w/o initialization). Tab. 3 shows that when removing the proposed initialization, there is a significant decrease in the image-text alignment of Story-Adapter in terms of CLIP-T. Fig. 4.4 illustrates that without initialization, the diffusion model fails to generate the required objects according to text prompts, e.g., “nightingale” and “robot”.
GRCA vs CSA.
We investigate GRCA and CSA in Tab. 3 and Fig. 4.4, using the outputs of the first iteration from Story-Adapter and StoryDiffusion, respectively. Though GRCA generates less visual consistency during the first iteration than CSA in terms of aCCS and aFID in Tab. 3, GRCA’s global comprehension improves the consistency of multiple characters throughout stories shown in Fig. 4.4. For example, GRCA effectively preserves the consistency of emerging characters (e.g., “the character 1900”) while CSA fails.
@@ 5 Conclusions and Discussions
We introduce Story-Adapter, an iterative framework that adapts pre-trained Stable Diffusion models for long story visualization. By using the generated images from previous iterations as references, our method maintains semantic consistency and enhances generative quality for fine-grained interactions throughout the story, effectively reducing error accumulation and avoiding the propagation of flaws. For efficiency, we propose a plug-and-play Global Reference Cross-Attention (GRCA) module, which utilizes global image embeddings to reduce computational costs while preserving essential image information flow. Extensive experiments demonstrate that Story-Adapter outperforms existing methods on the regular-length story visualization dataset, and shows strong results in long story visualization. These findings highlight the potential of our iterative paradigm to advance the quality and coherence of text-to-image story visualization.
@@ A.3 Iterative Paradigm
To address the aforementioned limitations, we propose an iterative paradigm in Story-Adapter (Fig. 9). We constantly consider all generated images in the previous iteration with an iterative mechanism and model on the global embeddings. Specifically, when generating for the $k_{th}$ image, we propose to implement Global Reference Cross-Attention (GRCA) on global embeddings from all generated images in the previous iteration.
@@ Discussion.
By using all generated images from the previous iteration as reference images to guide the current generation, we effectively maintain semantic consistency throughout the story. Moreover, all the generated images as references are updated through each iteration. Taken together, the iterative paradigm effectively avoids the influence of defects in some reference images.
Appendix B Subject-Consistent Generation Comparison

### v2

@@ 4 Experiments
4.1 Experimental Setting
We use the StorySalon dataset (Liu et al., 2024) to benchmark performance for regular-length story visualization. Compared to the closed Flintstones (Gupta et al., 2018) and Pororo (Kim et al., 2017) benchmarks, the open-ended StorySalon benchmark provides an evaluation of more diverse scenarios. For long story visualization, we curate multiple long stories using GPT-4o (OpenAI, 2024). To evaluate the efficacy of Story-Iter, we report CLIP text-image similarity (CLIP-T) (Radford et al., 2021), average Fréchet Inception Distance (aFID) (Cheng et al., 2024b), and Character-Character Similarity (aCCS) (Cheng et al., 2024b). CLIP-T is to measure image-text alignment, and both aFID and aCCS are used to evaluate semantic consistency among generated images. We use 10 iterations and the fixed random seed for Story-Iter in experiments by default. For story visualization within 100 frames, 10 iterations are sufficient to maintain the consistency of the story sequence. Iteration number ablation in Sec. F. Human evaluation and diversity can be found in Sec. C and Sec. H.3.
Table 1: Quantitative comparison for regular-length stories.
@@ 4.1 Experimental Setting
IP-Adapter (Ye et al., 2023) 0.241 0.737 25.18
Story-Iter (Ours) 0.305 0.760 16.52
IP-AdapterXL (Ye et al., 2023) 0.244 0.758 14.70
@@
Story-IterXL (Ours) 0.310 0.818 14.63
@@
Story-Iter (Ours) 0.307 0.754 98.51
FLUX.1.Kontext (Batifol et al., 2025) 0.290 0.783 78.03
Nano Banana 0.310 0.791 106.14
@@
DALL·E 3 (Betker et al., 2023) 0.301 0.674 251.71
SEED-LLaMA (Ge et al., 2023) 0.279 0.688 88.96
ConsiStory (Tewel et al., 2024) 0.316 0.761 108.83
1prompt1story (Liu et al., 2025) 0.316 0.790 99.25
@@
Story-IterXL (Ours) 0.318 0.802 94.30
4.2 Regular-length Story Visualization
Based on the standard setup on StorySalon dataset (Liu et al., 2024), we compare with the following baselines: StoryDiffusion (Zhou et al., 2024), StoryGen (Liu et al., 2024), AR-LDM (Pan et al., 2024), SDM (Rombach et al., 2022), Prompt-SDM (Rombach et al., 2022), Finetuned-SDM (fine-tuned on StorySalon) (Rombach et al., 2022), and IP-Adapter (individual image generation) (Ye et al., 2023). Autostudio (Cheng et al., 2024a) requires the additional bounding boxes as prompt, thus it is not included in comparison. For Prompt-SDM, we use prompts of “cartoon-style images”. For StoryDiffusion, we follow the official setup by using the initial four generated images as the reference. To adhere to copyright restrictions and ensure fair comparisons, we exclusively utilize text prompts from the open-source subset of the StorySalon test set for evaluation. This subset comprises 6,026 prompts, with an average of 14 frames per story and the longest story up to 44 frames.
Figure 8: Ablation study of the weighting strategies in the iterative paradigm: linear weighting strategy $\lambda_{i}=\lambda_{1}+q\times(i-1)$ for the $i_{th}$ iteration or fixed weighting strategy (e.g. $\lambda=0.3,\,\lambda=0.5$). Zoom in for a better view. See Sec. F for results with more iterations.
@@ 4.2 Regular-length Story Visualization
CLIP-T results in Tab. 2 show that Story-Iter and StoryDiffusion (Zhou et al., 2024) visualize content more aligned to the text prompt than previous story visualization models (AR-LDM (Pan et al., 2024) and StoryGen (Liu et al., 2024)). Since the story visualization framework introduces additional image conditioning, it inevitably affects text alignment compared to SDM, which is conditioned only on text. Meanwhile, since neither Story-Iter nor most baselines are trained on the StorySalon dataset, we introduce aFID and aCCS metrics for a fair evaluation of the character consistency among generated story images. aFID and aCCS in Tab. 2 illustrate that Story-Iter achieves higher semantic consistency of the generated images compared to most baselines.
@@
In Fig. 6, we provide the qualitative comparison results of the open-ended story visualization. Although AR-LDM (Pan et al., 2024) and StoryGen (Liu et al., 2024) generate coherent image sequences based on story prompts, the quality of the generated images degrades when story length increases due to the error accumulation issue of the AR paradigm. Results of StoryDiffusion (Zhou et al., 2024) and Story-Iter show satisfactory story visualization performance. However, StoryDiffusion cannot maintain consistency between certain subjects due to lacking global story comprehension (e.g., “cat” in Fig. 6). Additionally, since StoryDiffusion requires the first few generated images as references, the visualization results are affected by the reference image flaws (e.g., “closed-eye issue” in Fig. 6). In comparison, Story-Iter outperforms in regular-length story visualization by generating coherent image sequences. These findings highlight the effectiveness of GRCA, as it incorporates all reference images to ensure semantic consistency.
@@ 4.3 Long Story Visualization
To better evaluate generative quality for long story visualization (i.e., up to 100 frames (Fig. 32 and Fig. 33)), we compare to existing story visualization methods, reference-image based image editing models, and reference-guided image generation methods. We use GPT-4o to generate 20 long story cases of 10 50-sentence and 10 100-sentence narratives.
@@
The quantitative results in Tab. 2 show that Story-Iter significantly improves the semantic consistency and the generative coherence for fine-grained interactions for long story visualization compared to existing models (Pan et al., 2024; Liu et al., 2024; Betker et al., 2023; Ge et al., 2023; Tewel et al., 2024; Liu et al., 2025). A key difference between our Story-Iter and StoryDiffusion (Zhou et al., 2024) lies in the paradigm: by refining frame generation with global (visual) and local (text) semantic context in the iteration paradigm, Story-Iter achieves higher consistency compared to StoryDiffusion, which relies on the fixed reference images. IP-Adapter (Ye et al., 2023) employs the same single reference image that leads to less aFID. In contrast, referencing full-length frames in our paradigm, our method ensures both visual consistency and text alignment. Compared with Story-Iter, reference-image editing models suffer from overly strong identity constraints and their inability to handle newly introduced characters, which makes them unsuitable for long-story visualization. See Sec. D for more comparisons with reference-image editing models.
@@
Fig. 7 shows the visualization results for long stories, indicating that Story-Iter can generate high-quality, thematically consistent long image sequences based on the text prompts. The results of the IP-Adapter (Ye et al., 2023) exhibit significant missing or incorrect interactions due to limited text controllability. Our method outperforms IP-Adapter in global context visualization, benefiting from the proposed iterative paradigm and referencing full-length story-frames, in contrast to single-image guided generation. StoryGen (Liu et al., 2024) suffers from a more serious error accumulation issue from the AR paradigm during extended storytelling. Without modeling the holistic visual context, StoryDiffusion (Zhou et al., 2024) and ConsiStory (Tewel et al., 2024) fail to generate faithful character interactions and maintain consistency throughout the story. Our Story-Iter refers to the full-length story frames and individual text prompts in external iterations, thus achieving more accurate character interactions and maintaining subject consistency.
4.4 Variants
[…]
Method Diffusion Steps External Iterations FLOPs VRAM Time CLIP-T aCCS aFID
StoryDiffusion (Zhou et al., 2024) 50 1 22 PFLOPs 40GB 31 min 0.315 0.768 102.44
Story-IterXL 50 10 43 PFLOPs 19GB 250 min 0.318 0.802 94.30
Story-IterXL-Fast 4 10 3 PFLOPs 19GB 20 min 0.309 0.788 109.13
Story-Iter-Fast.
To improve the computational efficiency of Story-Iter for long-story visualization, we develop a fast variant built upon the SDXL-LCM backbone (Luo et al., 2023). This version reduces the diffusion sampling steps from 50 to only 4, substantially accelerating the iterative generation process while maintaining story-level consistency comparable to the original Story-Iter (see Tab. 3 and Sec. H.2).
Figure 9: Pose-conditioned generation results of Story-Iter-ControlNet. Given pose maps, the model produces frames that closely match the target poses while maintaining global story coherence.
Story-Iter-ControlNet.
We introduce a new variant, Story-Iter-ControlNet, which integrates a pose-conditioned ControlNet (Zhang et al., 2023) into the Story-Iter framework. This variant enables explicit control over character pose and spatial layout while fully preserving the global story consistency and text–image alignment achieved by Story-Iter. As shown in Fig. 9, by using pose maps (e.g., OpenPose skeletons) as control signals, Story-Iter-ControlNet enhances local-frame diversity without compromising global consistency or text–image alignment.
4.5 Ablation Study
Global Reference Cross-Attention.
We ablate the effect of global semantics modeling by GRCA for long story visualization. Specifically, for each image visualization in the sequence, we only use the single reference image at the corresponding index during the iteration as guidance. By establishing a global comprehension of the story for the diffusion model, Story-Iter maintains the semantic consistency in the generated image sequence (Tab. 4 and Fig. 10). For further discussion on GRCA, please refer to the Sec. A.5.
@@ 4.5 Ablation Study
Table 4: Ablation of the design choices of Story-Iter.
@@
Table 5: Effect of different weighting strategies in the Story-Iter. $\left[\lambda_{1},\lambda_{L}\right]$ denotes a linear weighting strategy with a lower bound of $\lambda_{1}$ and an upper term of $\lambda_{L}$ in 10 iterations.
fixed weighting strategy CLIP-T $\uparrow$ aCCS $\uparrow$ aFID $\downarrow$
$\lambda=0.3$ 0.320 0.760 101.55
$\lambda=0.5$ 0.261 0.753 81.72
linear weighting strategy CLIP-T $\uparrow$ aCCS $\uparrow$ aFID $\downarrow$
$\left[\lambda_{1},\lambda_{L}\right]=\left[0.2,0.5\right]$ 0.320 0.781 101.44
$\left[\lambda_{1},\lambda_{L}\right]=\left[0.3,0.4\right]$ 0.318 0.790 98.16
$\left[\lambda_{1},\lambda_{L}\right]=\left[0.3,0.5\right]$ 0.318 0.802 94.30
$\left[\lambda_{1},\lambda_{L}\right]=\left[0.3,0.6\right]$ 0.309 0.811 90.92
$\left[\lambda_{1},\lambda_{L}\right]=\left[0.4,0.5\right]$ 0.310 0.808 95.73
@@
We conduct ablation experiments to evaluate the effect of the proposed iterative paradigm for long story visualization. As shown in Tab. 4 and Fig. 8, the iterative paradigm improves generation quality for fine-grained interactions and semantic consistency. This is mainly because the iterative paradigm offers a global view of the entire story, thus reducing error accumulation and alleviating the propagation of the reference image flaws.
In Tab. 5 and Fig. 8, we validate our linear weighting strategy by comparing it with the fixed weighting strategy. While a smaller fixed weight factor enhances early text-image alignment, it limits consistency in later stages. Conversely, a larger fixed weight factor enforces excessive consistency across iterations, reducing flexibility. Thus, the fixed weighting strategy proves to be suboptimal. Based on the ablation results in Tab. 5, we adopt a linear weighting strategy: $\left[\lambda_{1},\lambda_{L}\right]=\left[0.3,0.5\right]$ in 10 iterations. This enables greater flexibility within the iterative paradigm.
@@
To ablate the effect of the initialization, we use a sequence of images consisting of the characters as reference images (i.e., w/o initialization). Adopting the same character as the reference image for the story sequence resulted in a high consistency of all story visualization results, triggering a better aFID. However, Tab. 4 shows that when removing the initialization, there is a significant decrease in the image-text alignment of Story-Iter in terms of CLIP-T. Fig. 10 illustrates that without initialization, the diffusion model fails to generate required objects according to texts.
@@ 5 Conclusions and Discussions
We introduce Story-Iter, a training-free iterative paradigm for long story visualization. By using full-length generated images from the previous external iteration as reference images, our framework maintains semantic consistency and enhances generative quality for fine-grained interactions throughout the story, effectively reducing error accumulation and avoiding the propagation of flaws. Extensive experiments demonstrate that Story-Iter outperforms existing methods on the regular-length story visualization dataset, and shows strong results in long story visualization. These indicate the potential of the proposed iterative paradigm to advance long story visualization.
6 Ethics Statement
All authors of this work have read and commit to adhering to the ICLR Code of Ethics.
7 Reproducibility
To ensure reproducibility, we provide pseudocode in Algo. 1 and implementation details in Sec. A.4. The full code can be found in the Supplementary Material.
@@ Discussion.
Algorithm 1 Pseudo-Code of Story-Iter.
[…]
# Initialize $I_{k,j}^{0}$, $I_{k,j}^{i}$~N(0, I), k~(1, B), i~(1, L), j~(0, J)
# Initialize Story-Iter iteration
for j in reversed(range(0, J)):
# Init z~N(0, I) if j>1 else z=0
[…]
R=concat($x_{1}^{0}$$x_{k}^{0}$$x_{B}^{0}$]), $x_{k}^{0}$=D($I_{k,0}^{0}$)
# Insert GRCA to $\theta$ and initialize weighting factor list $\lambda_{list}$
$\lambda_{list}$=linspace($\lambda_{s}$$\lambda_{e}$,L)
# Story-Iter Iteration
for $\lambda i$n enumerate($\lambda_{list}$):
for j in reversed(range(0, J)):
[…]
R=concat($x_{1}^{i}$$x_{k}^{i}$$x_{B}^{i}$]), $x_{k}^{i}$=D($I_{k,0}^{i}$)
Figure 12: Attention maps for each reference image in the ith iteration of Story-Iter for each image generation.
Table 6: GRCA vs. CSA.
Setting CLIP-T $\uparrow$ aCCS $\uparrow$ aFID $\downarrow$
GRCA 0.322 0.757 105.17
CSA 0.315 0.768 102.44
Setting FLOPs $\downarrow$ Time $\downarrow$ VRAM $\downarrow$
GRCA 44.09T 15S 19GB
CSA 22.32P 19S 40GB
Figure 13: Computational feasibility with scaling reference images.
@@ A.3 Iterative Paradigm
To address the aforementioned limitations, we propose an iterative paradigm in Story-Iter (Fig. 11). We constantly consider all generated images in the previous iteration with an iterative mechanism and model on the global embeddings. Specifically, when generating for the $k_{th}$ image, we propose to implement Global Reference Cross-Attention (GRCA) on global embeddings from all generated images in the previous iteration.
@@ Discussion.
By using all generated images from the previous iteration as reference images to guide the current generation, we effectively maintain semantic consistency throughout the story. Moreover, all the generated images as references are updated through each iteration. We illustrate in Fig. A.2 how different reference images contribute to the generation of each image within the iterative paradigm. Taken together, the iterative paradigm effectively avoids the influence of defects in some reference images. We demonstrate the procedure of Story-Iter in Algo. 1.
A.4 Implementation Details
To ensure a fair comparison, we used the weights of IP-Adapter (Ye et al., 2023) and IP-AdapterXL (Ye et al., 2023), respectively, resulting in two models: Story-Iter and Story-IterXL. We utilized DDIM (Song et al., 2020) for 50-step sampling with a Classifier-Free Guidance score set to 7.5. For the hyperparameters in our iterative paradigm, we set the number of story iterations to 10 by default. The weight factor $\lambda$ is set to 0.3 for the initial iteration and 0.5 for the final iteration, with linearly interpolated values for the intermediate iterations by our linear weighting strategy. Story-Iter is implemented in the NVIDIA RTX A6000.
Figure 14: Qualitative comparison of GRCA and CSA.
A.5 GRCA vs CSA
We investigate GRCA and CSA in Tab. 6 and Fig. 14, using the outputs of the first iteration from Story-Iter and StoryDiffusion, respectively. Though GRCA generates less visual consistency during the first iteration than CSA in terms of aCCS and aFID in Tab. 6, GRCA’s global comprehension improves the consistency of multiple characters throughout stories shown in Fig. 14. For example, GRCA effectively preserves the consistency of emerging characters (e.g., “the character 1900”), while CSA fails.
Establishing global story comprehension is fundamental to long-story visualization. Although the CSA can establish connections among all generated images during the generation process, its scalability for reference image scaling is limited, particularly for long stories, due to the high memory consumption of latent denoising representations. In Fig. 13, we analyze reference scaling of GRCA in single iteration and CSA in StoryDiffusion, without adopting the RI paradigm. FLOPs are calculated within the diffusion UNet. As shown in Fig. 13, as the number of reference images increases, StoryDiffusion experiences a significant rise in computation in terms of FLOPs, eventually becoming computationally prohibitive and can’t maintain global story comprehension of long narratives. While Story-Iter and Story-IterXL are slightly affected. This lays the foundation for GRCA to preserve global story semantics in long story visualizations.
[…]
A.6 GRCA Setting Ablation
Table 7: Ablation study on GRCA settings.
Setting CLIP-T aCCS aFID
Sliding Window 0.313 0.774 101.08
Random Reference Set 0.306 0.750 119.52
Mean-pooled Global Embedding 0.311 0.741 108.19
Story-Iter 0.318 0.802 94.30
To systematically evaluate how different configurations of GRCA affect the generation quality of Story-Iter, we conduct a series of ablation experiments by replacing GRCA with simpler global context aggregation settings (see Tab. 7). All variants are tested under the same iterative generation paradigm to isolate the contribution of long-range dependency modeling.
Sliding Window.
This variant aggregates only the most recent 10 frames, capturing short-range temporal dependencies but failing to model long-range narrative structure. When characters or objects reappear after long intervals, the model lacks sufficient global reference cues, leading to degraded long-story consistency.
Random Reference Set.
Random sampling introduces unstable and conflicting contextual signals. Such inconsistency accumulates across iterations and is particularly harmful in multi-character narratives, where mismatched global cues result in incoherent visual continuity.
Mean-pooled Global Embedding.
Averaging all global features compresses scene-specific and character-specific information into a single homogenized vector, causing significant loss of structural and semantic cues. As a result, the model struggles to preserve cross-scene consistency.
These results demonstrate that long-range, content-adaptive attention over the entire story sequence is essential, and simple aggregation methods cannot replace GRCA without significant quality degradation.
Appendix B Evaluation Metric
Figure 15: Calculation of aCCS and aFID.
We use Grounding-DINO (Liu et al., 2023) to extract object bounding boxes, using the name of each repeated object in the story as the grounding text. For each repeated object, the first frame in which it appears is treated as the reference frame. aCCS (Cheng et al., 2024b) is defined as the average CLIP-I similarity between the reference crop and its corresponding crops in subsequent frames. aFID (Cheng et al., 2024b) is defined as the FID between the reference crop and the distribution of later crops of the same object. Because repeated objects appear across multiple frames and all reference crops are obtained from text-conditioned generation, these metrics are stable and not overly sensitive to detection noise (see Fig. 15).
Appendix C Human Evaluation
Setting.
To complement the evaluation metrics to accurately reflect the quality of the generated stories, we involve human evaluation to further compare Story-Iter and baselines. Referring to the setting in StoryGen (Liu et al., 2024), we invite participants to rate various aspects: text-image alignment (Align.), character interaction (Inter.), content consistency (Cons.), and preference (Pref.) on a scale from 1 to 5. The higher the better.
Subject-Consistent Image Generation
Model Align. $\uparrow$ Inter. $\uparrow$ Cons. $\uparrow$ Pref. $\uparrow$
IP-Adapter (Ye et al., 2023) 2.51 3.27 4.58 4.19
IP-AdapterXL (Ye et al., 2023) 2.66 3.36 4.72 4.26
PhotoMaker (Li et al., 2024b) 3.79 4.18 4.25 4.11
StoryDiffusion (Zhou et al., 2024) 4.15 4.28 4.50 4.48
Story-Iter 4.02 4.20 4.41 4.33
Story-IterXL 4.20 4.35 4.58 4.54
Regular-Length Story Visualization
SDM (Rombach et al., 2022) 4.11 2.37 2.01 1.14
Prompt-SDM (Rombach et al., 2022) 4.03 3.49 1.99 1.26
Finetuned-SDM (Rombach et al., 2022) 3.35 3.82 2.15 1.60
AR-LDM (Pan et al., 2024) 3.08 3.64 2.90 2.05
StoryGen (Liu et al., 2024) 3.72 4.17 3.83 3.39
StoryDiffusion (Zhou et al., 2024) 3.96 4.48 4.52 4.37
Story-Iter 3.89 4.21 4.36 4.10
Story-IterXL 4.06 4.60 4.74 4.62
Long Story Visualization
AR-LDM (Pan et al., 2024) 3.30 3.68 3.42 3.27
StoryGen (Liu et al., 2024) 3.51 4.06 3.88 3.51
IP-Adapter (Ye et al., 2023) 3.79 4.27 4.30 4.06
IP-AdapterXL (Ye et al., 2023) 3.83 4.23 4.61 4.11
StoryDiffusion (Zhou et al., 2024) 4.16 4.30 4.53 4.35
Story-Iter 3.97 4.15 4.42 4.29
Story-Iter 4.35 4.47 4.70 4.65
[…]
Method Diffusion Steps External Iterations FLOPs VRAM Time CLIP-T aCCS aFID
Nano Banana - 1 - - 33 min 0.310 0.791 106.14
FLUX.1.Kontext (Batifol et al., 2025) 50 1 371 PFLOPs 68GB 48 min 0.290 0.783 78.03
Story-IterXL 50 10 43 PFLOPs 19GB 250 min 0.318 0.802 94.30
Story-IterXL-Fast 4 10 3 PFLOPs 19GB 20 min 0.309 0.788 109.13
Figure 16: Overly rigid identity constraints for Reference-image based image editing models.
Figure 17: New character missing for Reference-image based image editing models.
Results.
Tab. 8 shows that our Story-Iter receives more preference from the participants. It is worth noting that although IP-Adapter receives higher scores for consistency in the subject-consistent image generation task, Story-Iter is more favored in text-image alignment and generating character interactions. For regular-length and long story visualization, Story-Iter is more preferred compared to baselines in most evaluation aspects, especially visual consistency and capability to generate character interactions. This is aligned with the quantitative measurement.
Appendix D Comparison with Editing Models
In this section, we provide a further comparative analysis between Story-Iter and the latest reference-image based image editing models.
Qualitative Evaluation.
As shown in Tab. 9, Fig. 16 and Fig. 17, although reference-image based editing models can help maintain object consistency in story visualization, they impose overly rigid identity constraints compared with Story-Iter. Models such as Flux.1.Kontext (Batifol et al., 2025), and Nano-Banana are designed to preserve the appearance of the reference image as strictly as possible. While this behavior is desirable for image editing, it introduces a critical limitation in long-story visualization: the character remains nearly identical across frames, regardless of changes in action, pose, or narrative context. This severely reduces intra-story diversity and leads to lower CLIP-T scores. Reference-image based editing models also struggle to handle new characters or objects that do not appear in the reference image. Long stories frequently introduce new entities, for which models like Flux.1.Kontext and Nano-Banana cannot maintain cross-frame consistency. This results in degraded CLIP-T performance and unstable behavior in multi-character scenarios. This limitation is reflected in the lower CLIP-T of Flux.1.Kontext, as well as its difficulty in maintaining consistency when multiple characters are present. In contrast, Story-Iter is designed for open-ended, multi-character narratives rather than single-image editing. It employs a training-free, content-adaptive GRCA mechanism that maintains consistency across diverse scenes and newly introduced entities, without relying on a fixed reference image.
Efficiency Comparison.
[…]
Appendix E Subject-Consistent Generation Comparison
Related Work.
Subject consistency is critical for tasks such as story visualization and video generation. Recent advancements in subject-consistent image generation have focused on reducing computation while maintaining consistency. Early approaches like Gal et al. (2022); Ruiz et al. (2023) require extensive fine-tuning, prompting more efficient methods (Ryu, 2023; Han et al., 2023; Kumari et al., 2023; Yuan et al., 2023; Tewel et al., 2024; Cao et al., 2023; Xiao et al., 2024). Notable progress (Li et al., 2024b; Avrahami et al., 2024; Ye et al., 2023; Li et al., 2024a; Wei et al., 2023; Xu et al., 2024) includes IP-Adapter (Ye et al., 2023) with its decoupled cross-attention module. However, IP-Adapter is a single image generation model, which cannot be directly applied to story visualization that requires establishing connections for all frames.